capacity of capacitor is 4*10-12farad.convert this value in pf

Grade:9

2 Answers

Aarti Gupta
9 years ago
Capacity of capacitor is given as = 4*10-12F.
Value in pF(pico-Faraday) = ?
Solution –
Since,1pF = 10-12F
or, 10-12F = 1pF
thus,
4 * 10-12F = 4*10-12/10-12pF
or, 4pF
Adarsh
8 years ago
4*10-12*1012pF=4pF 4*10-12farad=